OpenStudy (roxy_martinez):

Two opposite sides of a square with length x are increased by 3 units each and the other two opposite sides are decreased by 3 units each. If the area of the resulting rectangle is 27 unit2, what would be the width of the rectangle? Hint: area of rectangle = length × width a)3 b)6 c)9 d)-3 e)-6

OpenStudy (welshfella):

New length = x + 3' and new width = x - 3 So the area of the rectangle Is (x+ 3)(x - 3) = 27

OpenStudy (welshfella):

Can you solve this for x?

OpenStudy (roxy_martinez):

but which of the answers is the correct then? a)3 b)6 c)9 d)-3 e)-6

OpenStudy (roxy_martinez):

oh I need to solve the x ?

OpenStudy (welshfella):

I can't tell you that. If you find x this will be the length of the square . the width of rectangle will be x - 3

OpenStudy (roxy_martinez):

well a solve x a that give me this x=6, -6 the what? it could be B or E

OpenStudy (welshfella):

That's the side of the square which must be 6 (side can't be negative)

OpenStudy (roxy_martinez):

ok then the correct answer is B thank YOU so much :)

OpenStudy (welshfella):

What was done to 6 to make width of the rectangle No the answer is not 6.@

OpenStudy (welshfella):

Read the first sentence of the question.